HUBBLE SPACE TELESCOPE DAILY REPORT #3076 PERIOD COVERED: 0000Z (UTC) 03/19/02 - 0000Z (UTC) 03/20/02 Daily Status Report as of 079/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Five Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.2 Completed Two Sets of STIS/CCD 8906 (Hot Pixel Annealing) The Space Telescope Imaging Spectrograph (CCD) was used to measure the effectiveness of the CCD hot pixel annealing process by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. In addition, CTE performance is examined by looking for traps in a low signal level flat. There were no reported problems. 1.3 Completed NCS 8967 (Start NICMOS Cooling System) The NICMOS Cooling System was configured to re-cool the NICMOS detectors, to determine the optimal science operating temperature for the NICMOS detectors, and to verify stable operation. The goals are as follows: 1} cool the NICMOS to the coldest possible temperature as specified by the NICMOS team {73K} with HST at a hot attitude and the NCC compressor operating near its maximum allowable speed. 2} measure NICMOS detector performance during the cool-down and steady state operation of the NCS. 3} demonstrate stability {+/-0.1K} of the NICMOS detector temperature at the optimal science operating temperature and 4} demonstrate repeatability {+/-0.1 K} of NICMOS detector temperature following changes from the optimal science operating temperature. All tests were successful. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 0 Successful: 0 Scheduled Re-acquisitions: 0 Successful: 0 2.2 FHST Updates: Scheduled: 0 Successful: 0 2.3 Operations Notes: The operations team continues to execute its responsibilities from the GSFC STOCC. At 079/0816Z, the HST coarse attitude error is 2.46 degrees. The NICMOS Cryo Cooler (NCC) continues operating nominally. SSA transmitter #1 was turned on for the interval 078/1201Z to 078/1219Z, using ROP IC-2. The SSR-1 EDAC error counter was cleared at 078/1307Z per ROP SR-1A. As directed by an operations request, Vehicle Disturbance Test #3 (Part 3) executed nominally between 078/171000Z and 078/234000Z. This portion of VDT #3 was executed in real-time to capture NCC surging disturbances at the 15+ hour mark since the transition to NCC PID Operate, at the maximum compressor speed during surging. This test was a 4-orbit test at +V3 sunpoint. During the test interval, the fundamental NCC surging frequency was observed to be ~1.6 Hz, and higher harmonics of the surging frequency were also observed at integer multiples of the surging frequency, as expected. During VDT #3 (Part 1), the fundamental surging frequency was observed at ~3.0 Hz (within an hour of the transition to PID Operate), and during the Vibration Emittance Ground Test, the fundamental surging frequency was observed at 3.5 Hz (within seconds of the transition to the maximum surging frequency). These results are all consistent with nominal operation of the NCC during cooldown, as the fundamental NCC surging disturbance frequency drops as the system temperature drops. As the NCS cools, the NCC surging frequency and its harmonics are sweeping through HST system modal frequencies, resulting in HST dynamic responses that allow identification of the various structural modes of the telescope. The modal properties of HST will be used to assess attitude control law stability margins. Using ROP 09A, the telemetry format was changed at 078/2326Z. The following operations notes were performed during this operational day: FGS HV Monitoring during the NCS activities Coarse Attitude Determination during NCS Fill Operations FHST 2 Mounting Bracket Limit Change and Monitoring Continue FGS HV Monitoring until ACR3 Activation 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW